The lenght of a rectangle is 8cm less than twice its width. Find the dimensions of the rectangle is the area is 120cm squared
Answer by checkley71(8403)

L=2W-8
W=W
(2W-8)W=120
2W^2-8W-120=0
2W^2-8W-120=0
2(W^2-4W-60)=0
2(W-10)(W+4)=0
W-10=0
W=10cm ANSWER FOR THE WIDTH.
L=2*10-8
L=20-8
L=12cm ANSWER FOR THE LENGTH.
PROOF.
10*12=120
120=120